OpenStudy (anonymous):

A local carnival costs $5 to enter and 50₵ per ride. What is the marginal cost of the 5th ride?

OpenStudy (anonymous):

I thought the MC would be .50? Was I wrong?

OpenStudy (anonymous):

this question is quiet tricky. the asked question is how much is the 5th ride? my answer: given; 5=entry 50=ride there are two answers. (1) if he/she didn't get out and continue riding until the fifth ride (2) if he/she got one ride every entry. solution for (1) 1 entry= 5 +.50(5) ; since 5 ride = 5+.5(5)=5+2.5=7.5; total payment solution for (2) entry = 5(5) = 25; since 5 entry ride = .5(5)= 2.5; since 5 ride =27.5;total payment